The market demand for Brand x has been estimated as:

Qdx=1,500−3Px−0.05M−2.5Py+7.5PzQxd=1,500−3Px−0.05M−2.5Py+7.5Pz

where PxPx is the price of Brand x, M per capita income, PyPy is the price of Brand y, and PzPz is the price of Brand z. Assume that Px=$2,M=$20,000,Py=$4,andPz=$4Px=$2,M=$20,000,Py=$4,andPz=$4.

Given the values specified in the problem, the market demand for Brand x is Select one: a. Qdx=1,500−(3∗2)−(0.05∗20,000)−(2.5∗4)+(7.5∗4)=474 b. Qdx=1,500−(3∗2)−(0.05∗20,000)−(2.5∗4)+(7.5∗4)=1,414 c. Qdx=1,500−(3∗2)−(0.05∗20,000)−(2.5∗4)+(7.5∗4)=514 Jump to...
Answer #1

Answer

The correct answer is (c) Qdx=1,500 − (3∗2) − (0.05∗20,000) − (2.5∗4)+(7.5∗4) = 514

Demand is given by:

Qdx = 1,500−3Px − 0.05M − 2.5Py + 7.5Pz

It is also Given that : Px = $2, M = $20,000, Py = $4 and Pz = $4

Hence, Using demand and given information we get :

Demand for brand x = Qdx = 1,500−3Px − 0.05M − 2.5Py + 7.5Pz

= 1,500−3*2 − 0.05*20,000 − 2.5*4 + 7.5*4

= 514

Hence, the correct answer is (c) Qdx=1,500 − (3∗2) − (0.05∗20,000) − (2.5∗4)+(7.5∗4) = 514